# Transfer Learning DA Domain Adaptation

### v1

@@ Abstract
Direct image-to-graph transformation is a challenging task that solves object detection and relationship prediction in a single model. Due to the complexity of this task, large training datasets are rare in many domains, which makes the training of large networks challenging. This data sparsity necessitates the establishment of pre-training strategies akin to the state-of-the-art in computer vision. In this work, we introduce a set of methods enabling cross-domain and cross-dimension transfer learning for image-to-graph transformers. We propose (1) a regularized edge sampling loss for sampling the optimal number of object relationships (edges) across domains, (2) a domain adaptation framework for image-to-graph transformers that aligns features from different domains, and (3) a simple projection function that allows us to pretrain 3D transformers on 2D input data. We demonstrate our method’s utility in cross-domain and cross-dimension experiments, where we pretrain our models on 2D satellite images before applying them to vastly different target domains in 2D and 3D. Our method consistently outperforms a series of baselines on challenging benchmarks, such as retinal or whole-brain vessel graph extraction.
Keywords:
Image-to-graph Inductive Transfer Learning Supervised Domain Adaptation
@@ 1 Introduction
Representing physical relationships via graph representations has proven to be an efficient, versatile, and expressive concept with vast utility in machine learning. Prominent examples are road network graphs [2], neuron representations and connections in the brain [39], blood vessels [8], and cell interactions [48]. The underlying physical structures are typically captured as a voxelized image without regarding semantic content. Hence, constructing graph representations from images (image-to-graph) is one of the critical challenges for unlocking the full potential in many real-world applications [23].
Figure 1: Conceptual overview of our framework. We use a transformer for single-stage image-to-graph extraction. Our three methodological contributions enable knowledge transfer between vastly different domains in 2D and 3D.
Traditionally, image-to-graph extraction involves a complex multi-stage process of segmentation, skeletonization, and graph pruning, which leads to inaccuracies and information loss at each step [14]. These limitations negatively impact prediction accuracy and reduce the application to downstream tasks that require more information from the underlying image (e.g., [38, 37]). Hence, there is a clear need for machine learning solutions that facilitate a direct image-to-graph extraction [20, 4]. Recently, vision transformers have been proposed for this task and showed superior performance to traditional multi-stage graph extraction approaches [36, 47]. However, these approaches only work on large sets of annotated 2D data and have not been shown to generalize to diverse 3D datasets where graph-level annotations are even less common than pixel annotations.
To address this challenge, we adopt and extend concepts from the field of inductive Transfer Learning (TL) (see taxonomy in [32] and Supplementary Figure 1), which has not been explored for image-to-graph extraction. Specifically, we assume the availability of large sets of image-graph pairs in some source domains (e.g. 2D satellite images of roads) and a small set of labeled data from the target domain (e.g. 3D microscopy image of vascular networks) that can both be used to train our image-to-graph model.
1.0.1 Our contribution.
Guided by the hypothesis that the underlying graph representations of physical networks have an inherently high similarity across domains, we introduce a set of methodological innovations for image-to-graph synthesis:
@@
    We introduce a novel edge sampling loss that regularizes the training procedure for relationship prediction, which allows us to pretrain our transformer on datasets that are vastly different in terms of the number of objects and relations (Sec. 3.1).
@@
    We propose a framework for 2D pretraining for 3D image-to-graph synthesis. Our framework introduces a projection function from the source representation to a space similar to the target domain (Sec. 3.4).
In extensive experiments on six datasets, we demonstrate the utility of our methods and outperform all baselines, validating our initial hypothesis. Our method leads to four main results: 1) we show compelling improvements in image-to-graph synthesis on existing datasets; 2) our method enables image-to-graph synthesis in previously unsolvable sparse data scenarios; 3) our method outperforms self-supervised pretraining by a margin and 4) our method bridges dimensions in pretraining, i.e. we solve the previously unsolved direct image-to-graph inference for complex 3D vessel images by pretraining on 2D road data.
Further ablation studies highlight that each of our methodological contributions addresses a specific weakness in the state-of-the-art.
@@ 2 Related Works
2.0.1 Image-to-graph extraction.
Image-to-graph extraction is an increasingly important field in computer vision with application to various domains, including road network extraction from satellite images [44] or 3D vessel graph extraction from microscopy images [3, 1]. Traditional methods solve these tasks through a multi-step approach that involves segmenting the input image [11, 2, 7], followed by skeletonization and pruning to generate the graph [44, 14]. Deep learning-based approaches frequently use an object detector followed by relation prediction [3, 1] or require a segmentation map as input [47]. The performance of these approaches is determined by the performance of the pipeline’s intermediate stages. Furthermore, information loss at each stage limits performance and applicability to downstream tasks. Lastly, these methods are tailored to specific domains, rendering them unsuitable for cross-domain applications, including TL.
@@
2.0.2 Transfer learning for transformers.
@@
Recent studies showed effective pretraining for transformers on natural images. Prevailing architectures involve supervised pretraining of the model’s backbone on, e.g., ImageNet, coupled with a random initialization of the encoder-decoder component [36, 29, 7]. Dai et al. [12] extend DETR [7] with a specific pretext task for object detection. Li et al. [25] compared self-and supervised pretraining methods with random weight initialization of a ViT-backbone [13]. They showed how self-supervised pretraining improves downstream object detection without a specific pretext task. A recent study [29] pretrained a transformer architecture on synthetic data and outperformed self-supervised methods for object detection. However, generating synthetic data, especially for specialized tasks, is typically domain-specific and requires expert knowledge [22]. Ma et al. [29] find that random weight initialization remains a robust baseline, often achieving comparable downstream performance to pretrained counterparts. To this date, no study has explored TL for image-to-graph transformers.
2.0.3 Cross-domain transfer learning.
@@
Existing cross-domain TL approaches are generative or discriminative. Generative approaches translate images from source to target domain on a pixel-level using a generative network [49]). In restricted settings, they have shown promising results, e.g., object detection for day-/night-time shifts in road scenes [21] or MRI to CT translation [19]. However, generative approaches require additional training of the translation network, which is computationally expensive and suffers from training instability [28]. Discriminative approaches train a single model to learn a general representation that is transferable between domains [43]. Some utilize a domain adversarial network that distinguishes whether a sample is from source or target domain based on its feature representation. Combined with a Gradient Reversal Layer (GRL) [16], this approach has proven effective in classification [16], segmentation [42], and object detection [10]. Chen et al. [10] introduced a Domain Adaptation (DA) framework for object detection using the GRL concept with two domain classifiers at image and instance level. Both are based on $\mathcal{H}$-divergence theory [5]. To reduce the bias towards the source domain, a consistency regularization penalizes the $L_{2}$-distance between the domain classifiers [10].
Still, to this day, all existing approaches are limited to a relatively small domain shift (e.g., from day to night scenes or synthetic to real images) instead of a fundamental domain shift, such as from satellite images to medical scans. Also, none of the existing approaches have been applied to the image-to-graph extraction problem studied in this work.
2.0.4 Cross-dimension transfer learning.
@@
2D to 3D TL is a highly challenging but promising research direction because of the abundance of labeled 2D data compared to the scarcity 3D data. The few existing approaches [35, 26, 46, 45] address the challenge, either data-based or model-based. Data-based approaches augment and project 2D data into 3D space, while model-based approaches aim to adjust the model to work with multi-dimensional input. Shen et al. [35] introduced a method that projects 3D point clouds to pseudo-2D RGB images. Liu et al. [26] introduced a pixel-to-point knowledge transfer that pretrains a 3D model by generating 3D point cloud data from 2D images with a learned projection function. A model-based approach by Xie et al. [46] used dimension-specific feature extractors and a dimension-independent Transformer. Similarly, Wang et al. [45] proposed a special tokenizer that creates 2D-shaped patch embeddings in a standard 2D ViT-model. All existing approaches either require changes to the target model, which may limit performance, a specifically crafted projection function, or additional training (e.g., for learning an explicit projection function). In our approach, we seek simplicity in implementation and training as well as generalizability to new domains and tasks.
@@ 3.1 Regularized Edge Sampling Loss
A leading difference between the source and target domain in our cross-domain TL setting is the different node and edge distribution, which poses a challenge because they dictate the relation loss calculation. Previous works arbitrarily pick the number of ground truth edges and fix it for the relation loss calculation ($\mathcal{L}_{\mathrm{rln}}$) for a specific dataset [36], which evidently generalizes poorly across varying domains. Generally, such a loss originates from object detection, where pair-wise relations are classified with a cross-entropy loss over a fixed number of edges $m$. This includes all active edges $\mathcal{A}$ and an irregular number of randomly sampled background edges $\bar{\mathcal{B}}\subseteq\mathcal{B}$. An active edge is defined as a pair of nodes ${g,h}$ that is connected by an edge, i.e., $e^{gh}_{rln}=1$. Similarly, for a background edge, $e^{gh}_{\mathrm{rln}}=0$ holds. Then, $\mathcal{L}_{\mathrm{rln}}$ is the cross-entropy loss on the set of sampled edges $\mathcal{R}=\mathcal{A}\cup\bar{\mathcal{B}}$ with $|\mathcal{A}|+|\bar{\mathcal{B}}|\leq m$:
@@
Then, the upsampled multiset of active edges $\mathcal{\hat{A}}$ is defined as:
@@
with: $|\mathcal{B}|\leq i<\frac{|\mathcal{E}|}{r}\quad\text{and}\quad i\in\mathbb{N}$. Note that effectively, only one of the sets is upsampled while the other stays the same because only one of the conditions in Eq. 5 or Eq. 6 can produce a valid $i$.
@@
3.2 Supervised Domain Adaptation for Transformers
In our setting, the stark differences between the source and target domain in image and graph features further amplify the TL challenge. Specifically, source and target domains significantly differ in image characteristics such as background and foreground intensities, signal-to-noise ratio, and background noise, as well as graph characteristics such as the structures’ radii or edge regularity. Edge regularity refers to the geometrical straightness of the underlying structure. While roads in the U.S. (e.g., highways) typically have a high edge regularity, vessels in microscopic images are highly irregular (i.e., the vessel does not follow a straight line and has high intra-edge curvature). To address this challenge, we utilize a domain adversarial on the image and graph level, respectively. These adversarial networks are used when jointly pretraining in both domains. Similar to previous methods [16, 41, 27], the image-level adversarial is a small neural network that classifies the domain based on a sample’s feature representation after the feature extractor. We treat each image patch $p^{u,v}$ at position $u,v$ as an individual sample and compute the cross-entropy loss
@@ 3.2 Supervised Domain Adaptation for Transformers
where $D\in[0,1]$ denotes whether the respective sample is from the source or target domain [10].
To align graph-level features, we view the concatenated tokenized transformer output $T\in\mathbb{R}^{(\#o+\#r)\times d}$ as a sample’s abstract graph representation where #$\#o$ and #$\#r$ are the numbers of object and relation tokens, respectively, and $d$ is the amount of the tokens’ hidden channels. We train a domain classifier on this abstract graph representation using the cross-entropy loss:
@@
Both domain classifiers are preceded by a GRL [16] reversing the gradient such that the main network is learning to maximize the domain loss. This framework forces the network to learn domain invariant representations and thus, aligns source and target domain. Furthermore, we apply a consistency regularization between both domain classifications to reduce the bias towards the source domain, as shown in [10]. This consistency regularization is expressed by an objective function minimizing the $L_{2}$-distance between the domain classifiers’ predictions (i.e., the output of the classification functions $d_{\mathrm{img}}$ and $d_{\mathrm{graph}}$):
@@ 3.3 Combined Training Loss
Our new regularized edge loss is combined with the other essential loss components to a final optimization function. The final loss consists of the $L_{1}$ regression loss ($\mathcal{L}_{\mathrm{reg}}$), the scale-invariant, generalized intersection over union loss ($\mathcal{L}_{\mathrm{gIoU}}$) for the box predictions (with predicted boxes $\tilde{v}_{box}$ and ground truth $v_{\mathrm{box}}$), and a cross-entropy classification loss ($\mathcal{L}_{\mathrm{cls}}$) for object classification [36]. Further, our new regularized edge sampling loss $\mathcal{L}_{\mathrm{Reslt}}$ and our three DA losses ($\mathcal{L}_{\mathrm{img}}$, $\mathcal{L}_{\mathrm{graph}}$, and $\mathcal{L}_{\mathrm{cst}}$) are included. Furthermore, in order to achieve unique predictions, we compute a bipartite matching between the ground truth and predicted objects utilizing the Hungarian algorithm [36]. $\mathcal{L}_{\mathrm{reg}}$, $\mathcal{L}_{\mathrm{gIoU}}$, and $\mathcal{L}_{\mathrm{Reslt}}$ are calculated over all object predictions $v$ that are matched to a ground truth, i.e. where $v_{\mathrm{cls}}^{i}=1$, whereas $\mathcal{L}_{\mathrm{cls}}$ is calculated over all object predictions. The combined loss term for our $N$ object tokens in a batch is defined as:
$\mathcal{L}=$ $\sum^{N}_{{i=1},\ {[v_{\mathrm{cls}}^{i}}=1]}\left[\lambda_{\mathrm{reg}}\mathcal{L}_{\mathrm{reg}}(v^{i}_{\mathrm{box}},\tilde{v}^{i}_{\mathrm{box}})+\lambda_{\mathrm{gIoU}}\mathcal{L}_{\mathrm{gIoU}}(v^{i}_{\mathrm{box}},\tilde{v}^{i}_{\mathrm{box}})\right]$ (11)
@@
$+\lambda_{\mathrm{DA}}(\mathcal{L}_{\mathrm{img}}+\mathcal{L}_{\mathrm{graph}}+\mathcal{L}_{\mathrm{cst}})+\lambda_{\mathrm{Reslt}}\underbrace{\sum_{\{g,h\}\in\hat{\mathcal{R}}}{\mathcal{L}_{\mathrm{CE}}(e^{gh}_{\mathrm{rln}},\tilde{e}^{gh}_{\mathrm{rln}})}}_{\mathcal{L}_{\mathrm{Reslt}}}$
with $\lambda_{\mathrm{reg}}$, $\lambda_{\mathrm{gIoU}}$, $\lambda_{\mathrm{cls}}$, $\lambda_{\mathrm{Reslt}}$, and $\lambda_{\mathrm{DA}}$ as respective weights.

### v2

@@ Abstract
Direct image-to-graph transformation is a challenging task that involves solving object detection and relationship prediction in a single model. Due to this task’s complexity, large training datasets are rare in many domains, making the training of deep-learning methods challenging. This data sparsity necessitates transfer learning strategies akin to the state-of-the-art in general computer vision. In this work, we introduce a set of methods enabling cross-domain and cross-dimension learning for image-to-graph transformers. We propose (1) a regularized edge sampling loss to effectively learn object relations in multiple domains with different numbers of edges, (2) a domain adaptation framework for image-to-graph transformers aligning image- and graph-level features from different domains, and (3) a projection function that allows using 2D data for training 3D transformers. We demonstrate our method’s utility in cross-domain and cross-dimension experiments, where we utilize labeled data from 2D road networks for simultaneous learning in vastly different target domains. Our method consistently outperforms standard transfer learning and self-supervised pretraining on challenging benchmarks, such as retinal or whole-brain vessel graph extraction. [^1] Code: github.com/AlexanderHBerger/cross-dim_i2g
@@ 1 Introduction
Representing physical relationships via graph representations has proven to be an efficient and versatile concept with vast utility in machine learning. Prominent examples are road network graphs [2], neuron representations and connections in the brain [45], blood vessels [8], and cell interactions [55]. Here, typically used voxelized images disregard the physical structure’s semantic content. Hence, constructing graph representations from images (image-to-graph, see Fig. 1) is a critical challenge for unlocking the full potential in many real-world applications [27].
Figure 1: Direct image-to-graph transformation. Whole brain vessel (top) and Agadez road dataset (bottom). The predicted graph is visualized as an overlay on the real image.
Figure 2: Conceptual overview of our framework. We use a transformer for single-stage image-to-graph transformation. Our three methodological contributions enable knowledge transfer between vastly different domains in 2D and 3D.
Traditionally, image-to-graph transformation involves a complex multi-stage process of segmentation, identification of physical structures and their relations, and iteratively pruning the constructed graph, which leads to inaccuracies and information loss at each step [14]. These disadvantages negatively impact prediction accuracy and limit the application to downstream tasks that require more information from the underlying image (e.g., [44, 43]). Hence, there is a clear need for machine learning solutions that facilitate accurate image-to-graph transformation in a single step [22, 4]. Recently, vision transformers have been proposed for this task and showed superior performance to traditional multi-stage approaches [42, 54]. However, these approaches require large sets of annotated 2D data and have not been shown to generalize to diverse 3D datasets where graph-level annotations are not widely available.
To address this challenge, we adopt and extend concepts from the field of inductive transfer learning (TL) (see taxonomy in [38]), which have not been explored for image-to-graph transformation. Inductive TL simultaneously uses large annotated datasets in the source domain (e.g., 2D satellite images) and a small set of labeled data in the target domain (e.g., 3D microscopy images of vascular networks).
Our contribution.
Guided by the hypothesis that the underlying graph representations of physical networks are similar across domains, we introduce a set of methodological innovations for image-to-graph transformation:
@@
    We introduce a loss formulation that regularizes the number of sampled edges during training, allowing simultaneous learning in domains that differ in terms of the number of objects and relations (Sec. 3.1).
@@
    We propose a framework for training 3D image-to-graph models with 2D data. Our framework introduces a projection function from the source representation to a space similar to the target domain (Sec. 3.4).
In extensive experiments on six datasets, we demonstrate our method’s utility and outperform all baselines, validating our initial hypothesis. Our method leads to four main results: 1) we show compelling improvements in image-to-graph transformation on existing datasets; 2) our method enables image-to-graph transformation in previously unsolvable sparse data scenarios; 3) our method outperforms self-supervised pretraining by a margin and 4) our method bridges dimensions during training, i.e. we solve the previously unsolved direct image-to-graph inference for complex 3D vessel images by utilizing 2D road data in an additional training step. Further ablation studies highlight that each of our methodological contributions addresses a specific weakness in the state-of-the-art.
@@ 2 Related Works
Image-to-graph transformation.
Image-to-graph transformation is an increasingly important field in computer vision with application to various domains, including road network extraction from satellite images [51] or 3D vessel graph extraction from microscopy images [3, 1]. Traditional methods solve this task through a multi-step approach involving input segmentation [11, 2, 7], skeletonization, and pruning to generate the graph [51, 14]. Deep learning-based approaches frequently use an object detector followed by relation prediction [3, 1] or require a segmentation map as input [54]. These approaches’ performances are determined by the performance of the pipeline’s intermediate stages. Furthermore, information loss at each stage limits performance and applicability to downstream tasks. Lastly, these methods are tailored to specific domains, rendering them unsuitable for cross-domain applications, including TL.
@@
Transfer learning for transformers.
@@
Recent studies showed effective pretraining for transformers on natural images. Prevailing architectures involve supervised pretraining of the model’s backbone on, e.g., ImageNet, coupled with random initialization of the encoder-decoder component [42, 33, 7]. Dai et al. [12] extend DETR [7] with a specific pretext task for object detection. Li et al. [28] compared self- and supervised pretraining methods with random initialization of a ViT-backbone [13]. They showed how self-supervised pretraining improves downstream object detection without a specific pretext task. Ma et al. [33] pretrained a transformer architecture on synthetic data and outperformed self-supervised methods for object detection. However, generating synthetic data, especially for specialized tasks, is typically domain-specific or requires expert knowledge [25]. Ma et al. [33] find that random weight initialization remains a robust baseline, often achieving competitive downstream performance. To this date, no study has explored TL for image-to-graph transformers.
Cross-domain transfer learning.
@@
Existing cross-domain TL approaches are generative or discriminative. Generative approaches translate images from source to target domain on a pixel-level using a generative network [56]. In restricted settings, they have shown promising results, e.g., object detection for day-/nighttime shifts in road scenes [23] or MRI to CT translation [21]. However, generative approaches require additional training of the translation network, which is computationally expensive and suffers from training instability [31]. Discriminative approaches train a single model to learn a general representation that is transferable between domains [50]. Some utilize a domain adversarial network that distinguishes whether a sample is from the source or target domain based on its feature representation. Combined with a gradient reversal layer (GRL) [17], this approach has proven effective in classification [17], segmentation [49], and object detection [10]. Chen et al. [10] introduced a domain adaptation (DA) framework for object detection using the GRL concept with two domain classifiers at the image and instance levels. Both are based on $\mathcal{H}$-divergence theory [5]. To reduce the bias towards the source domain, a consistency regularization penalizes the $L_{2}$-distance between the domain classifiers [10]. To this day, all existing approaches are limited to a relatively small domain shift (e.g., from day to night scenes or synthetic to real images) instead of a fundamental domain shift, such as from satellite images to medical scans. Also, none of the existing approaches have been applied to the image-to-graph transformation problem.
Cross-dimension transfer learning.
@@
2D to 3D TL is a highly challenging but promising research direction because of the abundance of labeled 2D data compared to the scarcity of 3D data. The few existing approaches [41, 29, 53, 52] address the challenge, either data-based or model-based. Data-based approaches augment and project 2D data into 3D space, while model-based approaches aim to adjust the model to work with multi-dimensional input. Shen et al. [41] introduced a method that projects 3D point clouds to pseudo-2D RGB images. Liu et al. [29] introduced a pixel-to-point knowledge transfer that pretrains a 3D model by generating 3D point cloud data from 2D images with a learned projection function. A model-based approach by Xie et al. [53] used dimension-specific feature extractors and a dimension-independent Transformer. Similarly, Wang et al. [52] proposed a special tokenizer that creates 2D-shaped patch embeddings in a standard 2D ViT-model. All existing approaches either require changes to the target model, which may limit performance, a specifically crafted projection function, or additional training (e.g., for learning an explicit projection function). In our approach, we seek simplicity in implementation and training as well as generalizability to new domains and tasks.
@@ 3.1 Regularized Edge Sampling Loss
A leading difference between the source and target domain in our cross-domain TL setting is the different node and edge distribution, which poses a challenge because they dictate the relation loss calculation. Previous works arbitrarily pick the number of ground truth edges and fix it for the relation loss calculation ($\mathcal{L}_{\mathrm{rln}}$) for a specific dataset [42], which evidently generalizes poorly across varying domains. Generally, such a loss originates from object detection, where pair-wise relations are classified with a cross-entropy (CE) loss over a fixed number of edges $m$. This includes all active edges $\mathcal{A}$ and an irregular number of randomly sampled background edges $\bar{\mathcal{B}}\subseteq\mathcal{B}$. An active edge is defined as a pair of nodes ${g,h}$ that is connected by an edge, i.e., $e^{gh}_{rln}=1$. Similarly, for a background edge, $e^{gh}_{\mathrm{rln}}=0$ holds. Then, $\mathcal{L}_{\mathrm{rln}}$ is the CE loss on the set of sampled edges $\mathcal{R}=\mathcal{A}\cup\bar{\mathcal{B}}$ with $|\mathcal{A}|+|\bar{\mathcal{B}}|\leq m$:
@@
Then, the upsampled multiset of active edges $\mathcal{\hat{A}}$ is:
@@
with: $|\mathcal{B}|\leq i<\frac{|\mathcal{E}|}{r}$. Notably, only one of the sets is upsampled while the other stays the same because only one of the conditions in Eq. 5 or Eq. 6 can produce a valid $i$.
@@
3.2 Supervised Domain Adaptation
In our setting, the stark differences between the source and target domain in image and graph features further amplify the TL challenge. Specifically, source and target domains significantly differ in image characteristics such as background and foreground intensities, signal-to-noise ratio, and background noise, as well as graph characteristics such as the structures’ radii or edge regularity. Edge regularity refers to the geometrical straightness of the underlying structure. While roads in the U.S. (e.g., highways) typically have a high edge regularity, vessels in microscopic images are highly irregular (i.e., the vessel does not follow a straight line and has high intra-edge curvature). To address this challenge, we utilize a domain adversarial network on the image and graph level, respectively. These adversarial networks are used when jointly pretraining in both domains. Similar to previous methods [17, 48, 30], the image-level adversarial network is a small neural network that classifies the domain based on a sample’s feature representation after the feature extractor. We treat each image patch $p^{u,v}$ at position $u,v$ as an individual sample and compute the CE loss
@@ 3.2 Supervised Domain Adaptation
where $D\in\{0,1\}$ denotes whether the respective sample is from the source or target domain [10].
To align graph-level features, we view the concatenated tokenized transformer output $T\in\mathbb{R}^{(\#o+\#r)\times d}$ as a sample’s abstract graph representation where #$\#o$ and #$\#r$ are the numbers of object and relation tokens, respectively, and $d$ is the amount of the tokens’ hidden channels. We train a domain classifier on this abstract graph representation using the CE loss:
@@
Both domain classifiers are preceded by a GRL [17] reversing the gradient such that the main network is learning to maximize the domain loss. This framework forces the network to learn domain invariant representations and, thus, aligns the source and target domain. Furthermore, we apply a consistency regularization between both domain classifications to reduce the bias towards the source domain, as shown in [10]. This consistency regularization is expressed by an objective function minimizing the $L_{2}$-distance between the domain classifiers’ predictions (i.e., the output of the classification functions $d_{\mathrm{img}}$ and $d_{\mathrm{graph}}$):
@@ 3.3 Combined Training Loss
Our new regularized edge loss is combined with the other essential loss components to a final optimization function. The final loss consists of the $L_{1}$ regression loss ($\mathcal{L}_{\mathrm{reg}}$), the scale-invariant, generalized intersection over union loss ($\mathcal{L}_{\mathrm{gIoU}}$) for the box predictions (with predicted boxes $\tilde{v}_{box}$ and ground truth $v_{\mathrm{box}}$), and a CE classification loss ($\mathcal{L}_{\mathrm{cls}}$) for object classification [42]. Further, our new regularized edge sampling loss $\mathcal{L}_{\mathrm{Reslt}}$ and our three DA losses ($\mathcal{L}_{\mathrm{img}}$, $\mathcal{L}_{\mathrm{graph}}$, and $\mathcal{L}_{\mathrm{cst}}$) are included. Furthermore, in order to achieve unique predictions, we compute a bipartite matching between the ground truth and predicted objects utilizing the Hungarian algorithm [42]. $\mathcal{L}_{\mathrm{reg}}$, $\mathcal{L}_{\mathrm{gIoU}}$, and $\mathcal{L}_{\mathrm{Reslt}}$ are calculated over all object predictions $v$ that are matched to a ground truth, i.e. where $v_{\mathrm{cls}}^{i}=1$, whereas $\mathcal{L}_{\mathrm{cls}}$ is calculated over all object predictions. The combined loss term for our $N$ object tokens in a batch is defined as:
$\mathcal{L}=$ $\sum^{N}_{{i=1},\ {[v_{\mathrm{cls}}^{i}}=1]}\Bigl{[}\lambda_{\mathrm{reg}}\mathcal{L}_{\mathrm{reg}}(v^{i}_{\mathrm{box}},\tilde{v}^{i}_{\mathrm{box}})$ (11)
$\quad\quad\quad\quad+\lambda_{\mathrm{gIoU}}\mathcal{L}_{\mathrm{gIoU}}(v^{i}_{\mathrm{box}},\tilde{v}^{i}_{\mathrm{box}})\Bigr{]}$
@@
$+\lambda_{\mathrm{DA}}(\mathcal{L}_{\mathrm{img}}+\mathcal{L}_{\mathrm{graph}}+\mathcal{L}_{\mathrm{cst}})$
$+\lambda_{\mathrm{Reslt}}\underbrace{\sum_{\{g,h\}\in\hat{\mathcal{R}}}{\mathcal{L}_{\mathrm{CE}}(e^{gh}_{\mathrm{rln}},\tilde{e}^{gh}_{\mathrm{rln}})}}_{\mathcal{L}_{\mathrm{Reslt}}}$
with $\lambda_{\mathrm{reg}}$, $\lambda_{\mathrm{gIoU}}$, $\lambda_{\mathrm{cls}}$, $\lambda_{\mathrm{Reslt}}$, and $\lambda_{\mathrm{DA}}$ as weights.